OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 16x + 74

OpenStudy (sirm3d):

transpose the constant 74 then add 64 to both sides of the equation

OpenStudy (anonymous):

how do i do that

OpenStudy (sirm3d):

\[y-74=x^2+16x\] \[y-74+64=x^2+16x+64\] \[y-10=(x+8)^2\]

OpenStudy (anonymous):

so it would be y= (x+8)^2 +10

OpenStudy (sirm3d):

yes

OpenStudy (anonymous):

thank you so much man you been so helpful